OpenStudy (anonymous):

find the 3rd, 4th, and 5th terms if: a1=3 a2=6 an=3an-1-2an-2

OpenStudy (anonymous):

The first term, a 1 = 3 The second term, a 2 = 6 I will show you how to find the 3rd term, a 3, and then you will find the 4th and 5th term. The third term, a 3 = ? So n = 3 (the third term) Your formula given is that

OpenStudy (anonymous):

a n = 3 a (n-1) - 2 a (n-2) So just substitute n = 3 , because that is what we want..so you get: a 3 = 3 a2 - 2 a 1 a 3 = 3(6) - 2(3) = 18 - 6 = 12 Now, you do the same for the fourth term, a 4. Plug in n = 4, just like I did for n = 3.